OpenStudy (anonymous):

If f (x)=4x and g (x)=x^2-2 what is g F (x)

OpenStudy (anonymous):

does this mean \[g\circ f(x)\]?

OpenStudy (anonymous):

yes thats what it means

OpenStudy (anonymous):

if so then \[g\circ f(x)=g(f(x))=g(4x)=(4x)^2-2=16x^2-2\]

OpenStudy (anonymous):

yes g(x) = (4x)^2 - 2 = 16x^2 - 2